OpenStudy (anonymous):

The difference of square of two numbers is 180. The square of the smaller number is 3 times the larger number. Find the two numbers.

OpenStudy (anonymous):

i got the answer but thats not the answer given in the book.

OpenStudy (helpsos):

I'm a little confused about the second sentence. The square of the smaller number is 3 times the larger number? So is this saying that when you multiply the larger number by 3 you get the square of the smaller number?

OpenStudy (anonymous):

it means that y^2=3x

OpenStudy (helpsos):

k

OpenStudy (kropot72):

Let the larger number = x \[x ^{2}-3x=180\] Solve to find x.

OpenStudy (kropot72):

Solve the quadratic\[x ^{2}-3x-180=0\]

OpenStudy (kropot72):

\[x ^{2}-3x-180=(x-15)(x+12)=0\]

OpenStudy (anonymous):

yes so one of the number is 15 right? but the answer says 12 and 6

OpenStudy (anonymous):

i think the answer given in book is wrong.

OpenStudy (anonymous):

the right answer is 15 and \[\sqrt{45}\]

OpenStudy (kropot72):

The difference of the two squares should equal 108 to give the answer in the book. I think that the question is wrong.

OpenStudy (nincompoop):

make sure it is not a typographical error

OpenStudy (kropot72):

If the question was to read as follows then the answer in the book is correct. The difference of square of two numbers is 108. The square of the smaller number is 3 times the larger number. Find the two numbers.

OpenStudy (anonymous):

i think that they took x=12 instead of -12. then they will get that answer